### v2

@@ 4.1.1 Disjoint SSGD
The results for online disjoint SSGD (i.e., SSep and VAD are trained separately) diarization models on Fisher are reported in Table 1. Results for CALLHOME are shown in Table 2. In this case, we do not fine-tune neither SSep nor VAD on CALLHOME data. Oracle sources refer to SSGD with oracle SSep, thus with error coming only from the VAD module. Note that oracle evaluation is missing for CALLHOME, as separated sources are not available. We also show the results of EEND on the CALLHOME test set, as reported in their respective original works.
@@
Table 3: End-to-end SSGD: separation and diarization results on the Fisher test set. The best results among proposed techniques are shown in bold.
Method Online Fair Eval. Full Eval. MS FA SC DER MS FA SC DER Disjoint SSGD w/LR 6.8 2.2 2.2 11.2 9.8 10.6 3.1 23.6 Disjoint SSGD w/LR (sim-CH adapt.) 6.7 2.2 0.6 9.5 6.5 11.2 1.5 22.1 VAD fine-tuning 6.8 2.3 0.6 9.8 10.9 6.9 1.3 19.1 VAD fine-tuning w/LR 7.6 1.5 0.8 9.8 12.5 5.3 1.5 19.3 SSep+VAD fine-tuning 6.3 2.1 0.4 8.8 8.8 7.0 0.7 16.5 SSep+VAD fine-tuning w/LR 6.5 1.9 0.4 8.8 9.3 6.5 0.8 16.6 Disjoint SSGD w/LR 6.2 2.6 1.5 10.2 8.5 12.1 2.3 22.9 Disjoint SSGD w/LR (sim-CH adapt.) 5.6 2.9 0.8 9.3 7.7 12.9 1.5 22.1 VAD fine-tuning 7.2 2.1 1.0 10.2 11.3 6.2 1.6 19.1 VAD fine-tuning w/LR 7.6 1.3 1.2 10.0 12.3 5.0 1.8 19.1 SSep+VAD fine-tuning 6.5 2.0 0.7 9.2 9.2 6.2 1.0 16.4 SSep+VAD fine-tuning w/LR 6.7 1.8 0.7 9.2 9.8 5.7 1.0 16.6
Table 4: End-to-end SSGD: diarization results on the CALLHOME test set. The best results among proposed techniques are shown in bold.
@@ 4.1.2 End-to-End SSGD
Method VAD Leakage Removal MS FA SC DER VBx (Landini et al., 2022) TCN n.a. 7.3 1.9 3.1 12.3 VBx (Landini et al., 2022) Kaldi 8.3 0.9 2.6 11.7      + Overlap assignment (Bullock et al., 2020) Kaldi 5.3 2.5 2.4 10.3 Spectral clustering (Park et al., 2019) Kaldi 8.3 0.9 5.3 14.5      + Overlap assignment (Raj et al., 2021b) Kaldi 5.7 2.7 5.8 14.1 SA-EEND (Fujita et al., 2019b) n.a. 9.5 SA-EEND-EDA (Horiguchi et al., 2020) n.a. 8.1 EEND + VC (Kinoshita et al., 2021b) n.a. 4.0 2.4 0.5 7.0 EEND-GLA (Horiguchi et al., 2021b) n.a. 6.9 DIVE (Zeghidour et al., 2021) n.a. 6.7 Conv-TasNet Energy 5.5 4.7 0.7 10.9 DPTNet 5.4 5.3 0.4 11.1 DPRNN 5.5 5.2 0.9 11.6 Conv-TasNet TCN 6.5 4.4 0.5 11.4 DPTNet 7.0 3.3 0.4 10.7 DPRNN 6.5 4.0 0.7 11.2 Conv-TasNet TCN 6.1 2.6 0.9 9.6 DPTNet 6.3 2.4 0.8 9.5 DPRNN 6.2 2.6 1.5 10.2
@@ 4.2 Offline Diarization
We compare the offline SSGD with both clustering-based and EEND methods. As clustering-based baselines we use VBx (Landini et al., 2022) and spectral clustering (Park et al., 2019), along with their overlap-aware counterparts (Bullock et al., 2020; Raj et al., 2021b). In this case, we use the publicly available Kaldi ASpIRE VAD  (Peddinti et al., 2015) model [^2] https://kaldi-asr.org/models/m4. For overlap detection, we fine-tune the Pyannote (Bredin et al., 2020) segmentation model [^3] https://huggingface.co/pyannote/segmentation on the full CALLHOME adaptation set. We tune the hyperparameters for each aforementioned module on the associated validation sets. To perform a fair comparison, we also tested the use of the TCN-based VAD in the VBx system, which however led to lower performance. For CALLHOME, we also report diarization errors of state-of-the-art EEND systems, as done in Table 2.
@@ 4.3 CSS Window Analysis
The results are shown in Fig. 2 for both datasets. Generally, we observed that the main source of error came from speaker confusion which consistently decreased for larger CSS windows, while missed speech and false alarm error rates remained approximately constant. In the presence of longer input recordings the CSS benefited from using longer windows to reduce errors in computing the correct permutation across frames during the stitching stage. Indeed, the correct permutation is computed on the overlapped portion using the cross-correlation, which is more reliable when computed on larger segments. The performances were close to the offline ones for windows larger than $60$ and $30$ s for Fisher and CALLHOME, respectively. The different slope of DER curves in Fig. 2(a) and 2(b) is due to the different average recording duration, which is $10$ minutes and $72$ s for Fisher and CALLHOME, respectively. This finding also suggests a parallelization strategy for offline CSS. Indeed, using shorter windows (e.g., $30$ or $60$ s) results in a lower memory footprint and higher inference speed-ups without affecting separation and diarization capabilities.
@@ 4.6 ASR Evaluation
Although the leakage removal post-processing generally improves diarization, the filled zeros could produce less natural utterances which negatively affect the WER. On the other hand, in the proposed framework it could be only exploited for obtaining the segmentation using the non-processed estimated sources (+ Leakage removal (seg-only)). In this latter case the DER is slightly reduced, as well for the VAD fine-tuning. Indeed, better segmentation improves performance when used with the same separated sources. The lower separation capability of models (cfr. Table 3) trained with the SSep+VAD fine-tuning resulted in higher WER as, during fine-tuning, there are no guarantees the output of the separator will be distortion-free, and even subtle distortions, while minimal in terms of SI-SDR, are known to affect ASR models significantly (von Neumann et al., 2020). Adding distortion-free constraints while fine-tuning, such as multi-frame minimum variance distortion-less response (Tammen and Doclo, 2021) is a possible future direction.
@@ 5 Conclusion and Future Work
Additionally, we have extended our previous work by performing a more comprehensive analysis of SSGD in real-world telephone conversation scenarios. We have experimented with another SSep model, i.e., DPTNet, which can be easily adapted to perform online inference with desired latency. Experimental results demonstrated that additional latency did not improve diarization accuracy, thus it was always convenient to use the model with the lowest latency. We have also generated a simulated dataset for the purpose of adaptation of SSep models to CALLHOME data. Adaptation on such data provided significant improvements, especially for DPRNN. The use of the proposed leakage removal algorithm was investigated both for disjoint and SSep+VAD fine-tuned SSGD. Experiments clearly showed that the post-processing algorithm was very effective to reduce both FA and DER in disjoint SSGD. However, when used in conjunction with fine-tuned models it did not improve the overall performance as the models learned to handle leaked segments directly from data. Finally, we have demonstrated that estimated sources generated by SSGD could be readily fed in input to an ASR system. The ASR performance largely improved over the input mixture and in some instance was even close to the one obtained with oracle sources. However, the end-to-end integration led to significant ASR performance degradation.
In future works, the SSGD framework will be extended to domains in which a higher number of speakers is typically involved (e.g., meeting scenarios). This will need likely the development of new techniques for speech separation. One of the major issues of having more than two speakers is that the speaker tracking for arbitrarily long inputs is significantly more difficult, potentially leading to confusion errors. Another issue is that the channel leakage problem also becomes more challenging as the number of output streams grows (e.g. $4$ channels for a max of $4$ local speakers). This latter can be addressed by adopting separation methods which can handle a variable number of output channels Nachmani et al. (2020). At the same time, novel strategies should be investigated to mitigate the separation/ASR degradation when using end-to-end fine-tuned SSGD models. This includes distortion-less constraints and/or continual learning approaches.
@@ 6 Acknowledgements
This work has been supported by the AGEVOLA project (SIME code 2019.0227), funded by the Fondazione CARITRO.

### v3

@@ 4.1.1 Disjoint SSGD
The results for online disjoint SSGD (i.e., SSep and VAD are trained separately) diarization models on Fisher are reported in Table 1. Results for CALLHOME are shown in Table 2. In this case, we do not fine-tune neither SSep nor VAD on CALLHOME data. Oracle sources refer to SSGD with oracle SSep, thus with error coming only from the VAD module. Note that oracle evaluation is missing for CALLHOME, as separated sources are not available. We also show the results of EEND on the CALLHOME test set, as reported in their respective original works. For all online EEND systems, with the exception of SA-EEND with STB (Xue et al., 2021b), the results are obtained estimating the number of speakers in input recordings as these systems are only tested in this setup.
@@
Table 3: End-to-end SSGD: separation and diarization results on the Fisher test set. The best results among proposed techniques are shown in bold. "w/LR" is appended to method name when leakage removal is applied.
Method Online Fair Eval. Full Eval. MS FA SC DER MS FA SC DER Disjoint SSGD w/LR 6.8 2.2 2.2 11.2 9.8 10.6 3.1 23.6 Disjoint SSGD w/LR w/Simu-CH adapt. 6.7 2.2 0.6 9.5 6.5 11.2 1.5 22.1 VAD fine-tuning 6.8 2.3 0.6 9.8 10.9 6.9 1.3 19.1 VAD fine-tuning w/LR 7.6 1.5 0.8 9.8 12.5 5.3 1.5 19.3 SSep+VAD fine-tuning 6.3 2.1 0.4 8.8 8.8 7.0 0.7 16.5 SSep+VAD fine-tuning w/LR 6.5 1.9 0.4 8.8 9.3 6.5 0.8 16.6 Disjoint SSGD w/LR 6.2 2.6 1.5 10.2 8.5 12.1 2.3 22.9 Disjoint SSGD w/LR w/Simu-CH adapt. 5.6 2.9 0.8 9.3 7.7 12.9 1.5 22.1 VAD fine-tuning 7.2 2.1 1.0 10.2 11.3 6.2 1.6 19.1 VAD fine-tuning w/LR 7.6 1.3 1.2 10.0 12.3 5.0 1.8 19.1 SSep+VAD fine-tuning 6.5 2.0 0.7 9.2 9.2 6.2 1.0 16.4 SSep+VAD fine-tuning w/LR 6.7 1.8 0.7 9.2 9.8 5.7 1.0 16.6
Table 4: End-to-end SSGD: diarization results on the CALLHOME test set. The best results among proposed techniques are shown in bold. "w/LR" is appended to method name when leakage removal is applied. "w/Simu-CH adapt." is appended when the separator is adapted on the simulated CALLHOME data.
@@ 4.1.2 End-to-End SSGD
Method VAD Leakage Removal MS FA SC DER VBx (Landini et al., 2022) TCN n.a. 7.3 1.9 3.1 12.3 VBx (Landini et al., 2022) Kaldi 8.3 0.9 2.6 11.7      + Overlap assignment (Bullock et al., 2020) Kaldi 5.3 2.5 2.4 10.3 Spectral clustering (Park et al., 2019) Kaldi 8.3 0.9 5.3 14.5      + Overlap assignment (Raj et al., 2021b) Kaldi 5.7 2.7 5.8 14.1 SA-EEND (Fujita et al., 2019b) n.a. 9.5 SA-EEND-EDA* (Horiguchi et al., 2020) n.a. 8.1 EEND-VC* (Kinoshita et al., 2021a) n.a. 4.0 2.4 0.5 7.0 EEND-GLA* (Horiguchi et al., 2021b) n.a. 6.9 DIVE (Zeghidour et al., 2021) n.a. 6.7 Conv-TasNet Energy 5.5 4.7 0.7 10.9 DPTNet 5.4 5.3 0.4 11.1 DPRNN 5.5 5.2 0.9 11.6 Conv-TasNet TCN 6.5 4.4 0.5 11.4 DPTNet 7.0 3.3 0.4 10.7 DPRNN 6.5 4.0 0.7 11.2 Conv-TasNet TCN 6.1 2.6 0.9 9.6 DPTNet 6.3 2.4 0.8 9.5 DPRNN 6.2 2.6 1.5 10.2
@@
*The number of speakers in input recordings is estimated by the model.
@@ 4.2 Offline Diarization
We compare the offline SSGD with both clustering-based and EEND methods. As clustering-based baselines we use VBx (Landini et al., 2022) and spectral clustering (Park et al., 2019), along with their overlap-aware counterparts (Bullock et al., 2020; Raj et al., 2021b). In this case, we use the publicly available Kaldi ASpIRE VAD  (Peddinti et al., 2015) model [^2] https://kaldi-asr.org/models/m4. For overlap detection, we fine-tune the Pyannote (Bredin et al., 2020) segmentation model [^3] https://huggingface.co/pyannote/segmentation on the full CALLHOME adaptation set. We tune the hyperparameters for each aforementioned module on the associated validation sets. To perform a fair comparison, we also tested the use of the TCN-based VAD in the VBx system, which however led to lower performance. For CALLHOME, we also report diarization errors of state-of-the-art EEND systems, as done in Table 2. For all the clustering-based baselines we assume that the oracle number of speakers (i.e., 2) is known in order to perform a fair comparison with the proposed SSGD approach. Regarding EEND approaches, for SA-EEND (Fujita et al., 2019b) and DIVE (Zeghidour et al., 2021) the maximum number of speakers is known, while for the other systems the speaker counting is estimated by the model. In particular, although the oracle number of speakers can be provided for EEND-VC, we decide to report the performance with estimated speaker counting as it leads to the best results in the 2-speaker scenario as reported in Kinoshita et al. (2021a).
@@ 4.3 CSS Window Analysis
(a) Fisher
(b) CALLHOME
The results are shown in Fig. 2 for both datasets Generally, we observed that the main source of error came from speaker confusion which consistently decreased for larger CSS windows, while missed speech and false alarm error rates remained approximately constant. In the presence of longer input recordings the CSS benefited from using longer windows to reduce errors in computing the correct permutation across frames during the stitching stage. Indeed, the correct permutation is computed on the overlapped portion using the cross-correlation, which is more reliable when computed on larger segments. The performances were close to the offline ones for windows larger than $60$ and $30$ s for Fisher and CALLHOME, respectively. The different slope of DER curves in Fig. 2(a) and 2(b) is due to the different average recording duration, which is $10$ minutes and $72$ s for Fisher and CALLHOME, respectively. This finding also suggests a parallelization strategy for offline CSS. Indeed, using shorter windows (e.g., $30$ or $60$ s) results in a lower memory footprint and higher inference speed-ups without affecting separation and diarization capabilities.
@@ 4.4 Latency Analysis with DPTNet
(a) Fisher
(b) CALLHOME
Figure 6: Diarization results of the disjoint SSGD with leakage removal on the test sets by varying leakage removal (LR) threshold $t_{\ell r}$.
@@ 4.5 Leakage Removal Analysis
Additionally, we investigate how varying the leakage removal threshold, $t_{\ell r}$, affects the DERs. The results are reported in Fig. 6. Both datasets showed similar trends. When $t_{\ell r}$ is very low the leakage removal algorithm greatly reduced false alarms introducing additional missed speech at the same time. On the other hand, for sufficiently high $t_{\ell r}$ values, the algorithm did not change the output and the metrics are on par with the one obtained by models without leakage removal. For all thresholds in between the overall performance was generally improved. The best DERs were obtained setting $t_{\ell r}$ to $-3$ and $3$, respectively. However, the best threshold for the DER metric highly degraded the SI-SDR on the Fisher test set. Then, we set the threshold $t_{\ell r}$ to $3$ for all experiments to achieve an optimal balance between separation and diarization capabilities.
@@ 4.6 ASR Evaluation
Although the leakage removal post-processing generally improves diarization, the filled zeros could produce fewer natural utterances which negatively affect the WER. On the other hand, in the proposed framework it could be only exploited for obtaining the segmentation using the non-processed estimated sources (+ Leakage removal (seg-only)). In this latter case the DER is slightly reduced, as well for the VAD fine-tuning. Indeed, better segmentation improves performance when used with the same separated sources. The lower separation capability of models (cfr. Table 3) trained with the SSep+VAD fine-tuning resulted in higher WER as, during fine-tuning, there are no guarantees the output of the separator will be distortion-free, and even subtle distortions, while minimal in terms of SI-SDR, are known to affect ASR models significantly (von Neumann et al., 2020). Adding distortion-free constraints while fine-tuning, such as multi-frame minimum variance distortion-less response (Tammen and Doclo, 2021) is a possible future direction.
4.7 Real Time Factor and Latency
To show that the proposed methods can be run in a truly online manner, we calculated the real time factor (RTF) of online DPRNN-based SSGD with and without the leakage removal module. The online processing unit length was of $0.1$ s, which corresponds to the algorithmic latency of online DPRNN-based SSGD. The RTF is computed as the ratio between the processing time and the length of an online processing unit. Since the end-to-end training strategies only change the model parameters without affecting the system architecture, the reported results are valid for both disjoint and end-to-end trained SSGD. We carried out our experiments on a Intel® Core i9-10920X CPU @ 3.50 GHz using one thread without any GPU. The RTF was equal to $0.159$ and $0.161$ for systems with and without leakage removal, respectively. As a result, the average latency time was about $0.116$ s for all DPRNN-based systems. This demonstrates that the proposed approach is applicable for real-time inference.
@@ 5 Conclusion and Future Work
Additionally, we have extended our previous work by performing a more comprehensive analysis of SSGD in real-world telephone conversation scenarios. We have experimented with another SSep model, i.e., DPTNet, which can be easily adapted to perform online inference with desired latency. Experimental results demonstrated that additional latency did not improve diarization accuracy, thus it was always convenient to use the model with the lowest latency. We have also generated a simulated dataset for the purpose of adaptation of SSep models to CALLHOME data. Adaptation on such data provided significant improvements, especially for DPRNN. The use of the proposed leakage removal algorithm was investigated both for disjoint and SSep+VAD fine-tuned SSGD. Experiments clearly showed that the post-processing algorithm was very effective to reduce both FA and DER in disjoint SSGD. However, when used in conjunction with fine-tuned models it did not improve the overall performance as the models learned to handle leaked segments directly from data. Finally, we have demonstrated that estimated sources generated by SSGD could be readily fed in input to an ASR system. The ASR performance largely improved over the input mixture and in some instances was even close to the one obtained with oracle sources. However, the end-to-end integration led to significant ASR performance degradation.
[…]
@@ 6 Acknowledgements
This work has been supported by the AGEVOLA project (SIME code 2019.0227), funded by the Trento and Rovereto Bank (CARITRO) Foundation.
Appendix A System Latency Computation
In this work, we use the expression “algorithmic latency” $l_{algo}$ to denote the latency obtained in an ideal setting in which processing time is equal to zero. It can be seen as the lower bound for a given algorithm. In particular, algorithmic latency is tied to length of online processing unit, which is the minimum amount of buffered data needed to produce new outputs. Latencies reported for all systems in Tables 1 and 2 follow this convention. In real-world setups, real latency $l_{real}$ also depends on hardware conditions. In this case, since the modules are connected in series, it is correct stating that we need to sum the algorithmic latency with the processing times of all modules (i.e., $t_{SSep}$, $t_{LR}$ and $t_{VAD}$):
$l_{real}=l_{algo}+t_{SSep}+t_{LR}+t_{VAD}\$ (2)
Below we provide a proof that the algorithmic latency of our system is equal to the largest latency among all modules. Fig. 7 shows a diagram of system latency computation in consecutive frames (which correspond to online processing units).
Figure 7: SSGD latency diagram.
In the online DPRNN-based SSGD, the speech separation (SSep) module has the largest algorithmic latency $l_{algo\_SSep}$. When the buffer is filled with an amount of speech equal to the SSep online processing unit length (which correspond to $l_{algo\_SSep}$), the SSep can process all the input data in the buffer. Since the algorithmic latency of the leakage removal (LR) module is lower than $l_{algo\_SSep}$, it can process all the SSep output when it is available. Then, only its processing time $t_{LR}$ contributes to the real latency $l_{real}$. The same applies to voice activity detection (VAD). Notice that during processing of frame $i-1$ the system can start to fill the buffer of next frame $i$. If the total processing time $t_{SSep}+t_{LR}+t_{VAD}$ is always lower than $l_{algo}$ (i.e., RTF $<1$) then $l_{real}$ represents the real latency for all frames. As a consequence, $l_{algo}=l_{algo\_{SSep}}$ and, in particular, $l_{real}=l_{algo}$ when total processing time is zero.